A company produces mosquito repellants that have a period of effectiveness that is approximately normally distributed with a mean of 300 hours and a standard deviation of 30 hours. A sample of 40 repellants was taken and the average period of effectiveness is computed as 290 hours. Is this sufficient evidence to conclude that the company’s claim is true at α=0.01?

1
Expert's answer
2022-05-31T11:26:23-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=300"

"H_1:\mu\not=300"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c =2.5758."

The rejection region for this two-tailed test is "R = \{z:|z|> 2.5758\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{290-300}{30/\sqrt{40}}=-2.1082"


6. Since it is observed that "|z|=2.1082<2.5758=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed is "p=2P(Z<-2.1082)=0.035014," and since "p=0.035014>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 300, at the "\alpha = 0.01" significance level.